OpenStudy (anonymous):

what is the distance between (-5,8) and (-3,6)???

OpenStudy (anonymous):

distance formula \[d = \sqrt{(x2 - x1)^2 + (y2-y1)^2}\]

OpenStudy (triciaal):

@Maddiestone119 please read the question before answering! first it asks for the distance between the 2 points not the slope second to find the slope its the rise /run and the rise is the change in the y value while the run is the change in the x value the slope using points (-5,8) and (-3,6) should be (6 -8) / (-3 - -5) = -2/2 = -1 third the slope is a ratio not a point @neb hope you will learn from this too. to find the distance use the distance formula given above by @hilbertboy96